OpenStudy (anonymous):

what is the answer/ i got 25/9 but need to know if its correct 2 1/3 - 1 1/2 * 3 1/3 =

OpenStudy (anonymous):

\[2\frac{1}{3}-1\frac{1}{2}*3\frac{1}{3}=\frac{7}{3}-\frac{3}{2}*\frac{10}{3}\] \[=\frac{7}{3}-\frac{30}{6}=\frac{14}{6}-\frac{30}{6}=\frac{-16}{6}=\frac{-8}{3}\]

OpenStudy (anonymous):

^^^ yes

OpenStudy (anonymous):

thanks, 1 more question i need help with

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

(3 1/4 - 1 2/3) * (5 1/3 - 2 1/2) =

OpenStudy (anonymous):

Try it out and show your work and I'll confirm it for you. Or guide you if its wrong.

OpenStudy (anonymous):

(13/4 - 5/3) (16/3 - 5/2) 39/12 - 20/12 32/6 - 15/6 19/12 * 17/6 323/72

OpenStudy (anonymous):

umm you wouldn't multiply the denominator it would stay as 12 that's the LCD

OpenStudy (anonymous):

idky it says i'm writing a reply when i'm not...

OpenStudy (anonymous):

Yes that is the correct answer, well done.

OpenStudy (anonymous):

\[((3+1/4)-(1+2/3))*((5+1/3)-(2+1/2))= \]\[\left(\frac{13}{4}-\frac{5}{3}\right)*\left(\frac{16}{3}-\frac{5}{2}\right)=\left(\frac{19}{12}\right)*\left(\frac{17}{6}\right)=\frac{323}{72} \]

OpenStudy (anonymous):

it's 323/12 why would the denominator change?

OpenStudy (anonymous):

the 12 is multiplied by the 6. \[(\frac{19}{12})(\frac{17}{6})\] \[\frac{19*17}{12*6}\] When multiplying fractions, you have to multiply through both numerator and denominator

OpenStudy (anonymous):

since when...lol